### v1

@@ 5.3.1 On the Model-based Controllers
The emergence of computationally efficient and accurate models enables researchers to develop model-based controllers that can accurately predict and react to the nonlinear nature of the elastic rods. As shown in Table 5, the controllers based on Inverse Kinematics are the most popular choice, achieving tip trajectory tracking tasks in real-world prototypes. Despite the simplified model, the feedback action is able to reject the unmodelled effects, showing the usability of that strategy.
@@
However, a dynamic model is essential to exploit the properties of continuum robots. Most of the proposed controllers adopt well-known approaches from the under-actuation literature, such as FL, ESC, SMC, or MPC. However, only a few proposed controllers presented an experimental validation. The reasons behind this sim-to-real gap could be the lack of robustness or physical limits of the prototypes. While the Inverse Dynamics, FL, and ESC are sensitive to unmodelled effects and parametric uncertainties, the discontinuous output of the SMC is difficult to implement with the most popular actuation technologies (e.g., tendons, pneumatic). Specifically, the actuators’ internal dynamics may limit the effective bandwidth of the system, hence filtering the discontinuous output of the controller. The MPC is a promising but not widely explored controller in the continuum robots literature.
Another relevant aspect of the analyzed works is the discretization technique employed. Two main aspects eased the control design: (i) the advancement of discretization techniques and (ii) the tunability of the accuracy. The former allows the researchers to design the controller with a finite number of state variables and compute efficiently the robot’s kinematics or dynamics. Selecting appropriate discretization techniques is crucial, particularly for determining the minimum frequency necessary to ensure numerical stability. In addition, the DER technique takes into account also cross-section deformations, especially useful for modeling and controlling muscles in the case of bio-inspired robotic systems (Chang et al., 2023). The latter aspect allows for the choice of accuracy levels and the number of strain modes the model can predict. A common approach simplifies the model and designs a controller that treats unmodeled strains as external disturbances. This strategy can enhance the speed of the control loop and its robustness.
@@
Regarding research gaps, it is possible to note the lack of experimental validation for MB controllers. Furthermore, there are few studies involving external forces due to interactions with environments, particularly because of the complexity involved in modeling and estimating contact and friction forces. Further developments on model-based controllers could benefit from using physics-inspired ANN to lessen the computational burden of complex expressions (e.g., the Coriolis matrix) or use them to compute the EoM of the elastic rod, as already explored in (Lutter and Peters, 2023; Liu et al., 2024). Finally, the literature on rod-based planning (Mishra et al., 2023; Bentley et al., 2023) is sparse, likely due to the current maturity of the controllers. Recently, some planning algorithms using only geometrical approaches have been proposed (Rao et al., 2024).
5.3.2 On the Learning-based Controllers
@@
Conversely, the role of rod models is central for developing complex deep RL control policies, especially to achieve unstructured physical interaction with CSM. Deep RL agents exploited computationally feasible rod models embedded in simulated environments to learn optimal iteratively. Herein, RL not only replaces the burden of collecting and labeling datasets offline but also unlocks a new range of possibilities not fully explored yet. The literature review highlighted the role of rod models in RL policies. Indeed, as reported in Table 6, control tasks evolved from position control for quasi-static tracking to simulated pose control for obstacle maneuvering and pose/force control for dynamic pushing. For instance, initial policies limited the physical interaction to static payloads without addressing the sim-to-real gap (Satheeshbabu et al., 2020). On the other hand, efforts in developing simulated environments (Naughton et al., 2021) combined with careful reward engineering and sim-to-real techniques (Alessi et al., 2024) facilitated recent advancements in physical interaction with CSM. In summary, rod models combined with recent deep RL methods and sim-to-real transfer techniques (Zhao et al., 2020b) advanced the state of the art of soft robot control.
Future developments could explore model-based RL. Here, the term model-based refers to the agent not only learning a policy but also learning a predictive model of the environment used for planning (Moerland et al., 2023). Alternative research lines can be borrowed from rigid robotics and tailored to the nonlinearities of continuum and soft robots (Ibarz et al., 2021).
@@ 6 Conclusion
This paper provided an overview of the modeling and control of continuum and soft robots using rod theories. First, we formulated the four main rod theories used in robotics under a unified mathematical framework. Subsequently, we provided mathematical tools for modeling the actuation in soft robotics. This was followed by a classification and summary of the rod-based models for continuum and soft robots across nine classes of deformations, offering a unique perspective on the evolution of the state of the art, together with a discussion on research trends and gaps that emerged. Furthermore, we reviewed recent control strategies leveraging rod models to perform various tasks, followed by a discussion on the encountered approaches as well as our insights. Through an in-depth exploration of various rod-based models and control strategies, the diverse range of applications of rod theories demonstrates their capabilities to capture the complex mechanical behaviors of continuum and soft robots. Despite the remarkable advancements in incorporating rod theories into soft robot models and controllers, there are still areas requiring further research. For instance, accurate modeling of processes inherent in soft materials, such as hysteresis, is still a challenge. Related to the control, addressing efficient sim-to-real policy transfer in learning-based approaches is essential for real-world robot deployment.
This work was supported by the Future and Emerging Technologies (FET) program of the European Commission, through the PROBOSCIS project (G.A. ID 863212), by the US Office of Naval Research Global under Grant N62909-21-1-2033, and in part by Khalifa University under Awards No. RIG-2023-048, RC1-2018-KUCARS

### v2

@@ 9.6 On the Model-based Controllers
The emergence of computationally efficient and accurate models enables researchers to develop model-based controllers that can accurately predict and react to nonlinear elastic rods. We observed that controllers based on IK are the most popular choice, as they achieve tip trajectory tracking tasks in real-world prototypes (Table 3). Despite the simplified model, the feedback action can reject unmodeled effects, showing the effectiveness of that strategy. However, we should note that most IK-based controllers treat the system as fully actuated. This limits control to a smaller set of DoFs than the robot is inherently capable of exhibiting.
@@
A dynamic model is essential to exploit the properties of continuum robots. Most of the proposed controllers adopt well-known approaches from the under-actuation literature, such as FL, ESC, SMC, or MPC. However, only a few proposed controllers were validated experimentally. The reasons for neglecting the sim-to-real comparison could be the lack of robustness or physical limits of the prototypes. While the Inverse Dynamics, FL, and ESC are sensitive to unmodelled effects and parametric uncertainties, the discontinuous output of the SMC is difficult to implement with the most popular actuation technologies (e.g., tendons, pneumatic). Specifically, the actuators’ internal dynamics may limit the effective system bandwidth, hence filtering the discontinuous output of the controller. The MPC is a promising but not widely explored controller in the continuum robots literature, which could be attributed to the relatively high computational cost.
Two key aspects facilitate the design of controllers for CSM: (i) the development of advanced discretization techniques and (ii) the tunability of model accuracy. The former enables the design of controllers with a finite number of state variables and allows for efficient computation of the robot’s kinematics and dynamics. Choosing appropriate discretization techniques is particularly critical for determining the minimum control frequency required to ensure numerical stability. Furthermore, the DER technique captures cross-sectional deformations, which are especially relevant for modeling and controlling muscle-like structures in bio-inspired robotic systems (chang2023energy). This feature was recently integrated within the GVS framework by sun2025real, who proposed an extended CRT formulation that models radial (in-plane) deformations through an additional strain variable. Their model also incorporates nonlinear hyperelastic and viscous effects, extending the GVS approach to more complex, bio-inspired dynamic scenarios. Equally important is the ability to tune model accuracy by neglecting specific modes or reducing the number of DoFs, as implemented in the Strain Parameterization. This flexibility allows trade-offs between accuracy and computational complexity while enabling the controller to treat unmodeled dynamics as external disturbances. Such a strategy increases the control loop frequency by reducing computational load and enhances its overall robustness.
@@
Regarding research gaps, we observed the lack of experimental validation for MB controllers. Furthermore, there are few studies involving external forces due to interactions with the environment, particularly because of the complexity involved in modeling and estimating contact and friction forces. Due to the current lack of maturity of the controllers, the literature on rod-based planning (mishra2023trajectory; bentley2023safer) is still sparse, although planning algorithms using geometrical approaches were recently proposed (rao2024towards). Future developments on model-based controllers could benefit from using physics-informed ANN to lessen the computational burden of complex expressions (e.g., the Coriolis matrix) or use them to compute the EoM of the elastic rod, as already explored in (lutter2023combining; liu2024physics). Another promising future direction is the application of the Collocated Form proposed in (pustina2024input) to the design of more advanced controllers, particularly those inspired by the rigid underactuated robotics literature. This specific change of coordinates can simplify feedback implementation, since the actuator lengths are directly measurable and easier to estimate, as also exploited by feliu2024dynamic. Furthermore, adopting this coordinate transformation can enhance IK-based controllers by naturally accounting for the system’s under-actuation. Finally, differentiable simulators and libraries for rod models could be utilized for optimal control. Recently, analytical derivatives for the GVS model were calculated in (mathew2024analytical), facilitating the implementation of model-based optimal controllers.
9.7 On the Learning-based Controllers
@@
Conversely, the role of rod models is central for developing complex deep RL control policies to achieve unstructured physical interaction with CSM. Deep RL agents exploited computationally feasible rod models embedded in simulated environments to learn optimal actions iteratively. Herein, RL reduces the burden of collecting labeled datasets offline and unlocks new possibilities that are not fully explored yet. The control tasks addressed evolved from position control for quasi-static tracking to simulated pose control for obstacle maneuvering and pose/force control for dynamic pushing (Table 4). For instance, initial policies limited the physical interaction to static payloads without addressing the sim-to-real gap (satheeshbabu2020continuous). Conversely, efforts in developing simulated environments (naughton2021elastica) combined with careful reward engineering and sim-to-real techniques (alessi2024pushing) facilitated recent advancements in physical interaction with CSM. In summary, rod models combined with recent deep RL methods and sim-to-real transfer techniques are promising to advance the manipulation capability of continuum soft robots.
We should note, however, that these works relied on motion capture systems to track the robot positions. Indeed, reliable soft robotic proprioception is still a challenge (pagliarani2025softtex). Future developments could explore model-based RL, where model-based here refers to the agent not only learning a policy but also learning a predictive model of the environment used for planning (moerland2023model). Alternative research lines can be borrowed from rigid robotics and tailored to the nonlinearities of continuum and soft robots (ibarz2021train).
@@ 10 Conclusion
This review paper explored the modeling and control of continuum and soft robots with a focus on rod theories. We surveyed the state of the art with a vertical approach spanning mathematical formulations of rod theories, rod-based models of continuum and soft robots, and control strategies with rod models. The review of the mathematical background facilitated comparison between different variants of rod theories while highlighting connections to other models. A comprehensive review of rod-based models uncovered their versatility in supporting various studies, from bioengineering design to experimental validation of continuum and soft robots. Meanwhile, our literature classification on deformation classes offered new perspectives on the evolution of the state of the art, especially in relation to robot deformations and the model. The review of controllers highlighted the direct applicability of advanced model-based algorithms for trajectory tracking tasks, while RL coupled with sim-to-real strategies emerged as a promising approach for manipulation.
Despite remarkable advancements in integrating rod theories within soft robot models and controllers, limitations persist. For instance, modeling hysteresis effects of soft materials and comprehensive experimental validations in contact scenarios remain challenging. Related to the control, addressing efficient sim-to-real policy transfer is essential for robot deployment. This paper offered a comprehensive treatment of rod models in continuum and soft robot control. Nonetheless, we believe our insights can benefit various soft robotics communities and could serve roboticists dealing with continuum or soft structures.
This work was supported by the Future and Emerging Technologies (FET) program of the European Commission, through the PROBOSCIS project (G.A. ID 863212), by the US Office of Naval Research Global under Grant N62909-21-1-2033, and in part by Khalifa University under Awards No. RIG-2023-048, RC1-2018-KUCARS, by the Singapore-Italy Cooperation Project DESTRO (A*STAR grant number R22I0IR124, _______ MAECI), by the MOE Tier 2 Project REBOT (grant number MOE-T2EP50221-0010), by the NUS grant RoboLife, and by the PNRR MUR project PE0000013-FAIR.